OpenStudy (anonymous):

Carrots sell for $2.10 per pound, and crackers sell for $2.90 per pound. Glen bought some carrots and some crackers. The total weight was 2.3 pounds and cost $6.03. How many pounds of carrots and how many pounds of crackers did Glen buy? Note: Please give a direct answer

OpenStudy (anonymous):

Probably one of theses answers. A. 0.5 pounds of carrots; 1.8 pounds of crackers B. 0.8 pounds of carrots; 1.5 pounds of crackers C. 1 pound of carrots; 1.3 pounds of crackers D. 1.5 pounds of carrots; 0.8 pound of crackers

OpenStudy (anonymous):

I will be more than happy to give you a direct explanation. First, you must use the information to create 2 equations you solve those by any method of elimination in order to get the values needed. I would suggest you choose some variables, I did A = carrots and R = crackers. Can you write the 2 equations?

OpenStudy (anonymous):

a + r = 6.03

OpenStudy (anonymous):

cost was 6.03 so that should be 2.10a + 2.90r = 6.03 pounds are 2.3 so that should be a + r = 2.3

OpenStudy (anonymous):

it's c or d

OpenStudy (anonymous):

by solving the equations we can get our answer. a + r = 2.3 2.10a + 2.90r = 6.03 multiply first equation by -2.10 (each term) now you can eliminate A and solve for R which tells you your pounds of crackers :)